OpenStudy (anonymous):

If the area of a circle is 36pi what is the circumference of the circle?

OpenStudy (nubeer):

12pi

OpenStudy (vishweshshrimali5):

2* pi * r = 2 pi 6 = 12 pi

OpenStudy (anonymous):

well, i the formula for the area of a circle is pi*r^2, this means that we have, in thiscase, the radius being 6, since 6 squared is 36. Now simplu plug the radius of 6 into the formula for the circufuremce of the circle, namely 2*pi*r

OpenStudy (anonymous):

\[pi * r^2=36 pi\] \[r^2=36\] \[r=6\] \[2*pi*r=2*pi*6=12 pi\]

OpenStudy (anonymous):

12*pi

OpenStudy (diyadiya):

area=pi * r^2 36pi = pi * r^2 r=6 2*pi*r=12pi